OpenStudy (anonymous):

Medal!!:) Choose the correct simplification of the expression f to the 6th power times h to the 9th power all over f to the 3rd power times h to the 7th power. f12h6 1 over f to the 12th power times h to the 6th power f3h2 1 over f to the 6th power times h to the 6th power 9. [7.04] Choose the correct simplification of the expression the quantity a to the 3rd power over b to the 7th power all raised to the 2nd power. a to the 5th power over b to the 9th power a over b to the 14th power a to the 6th power over b to the 14th power a6b14

OpenStudy (kamibug):

When dividing terms with exponents, the exponents are subtracted. So f^6 - f^3 = f^3 and h^9 - h^7 = h^2 So the answer to the first one is f^3 times h^2.

OpenStudy (kamibug):

I think the second one is D because when you raise a power to a power, you multiply the values so 3*2 = 6 and 7*2=14.